OpenStudy (anonymous):

sec theta=11/4, cot theta < 0

OpenStudy (anonymous):

1/cos x = 11/4 -> cos x = 4/11 = cos 68.68 deg and cos x = cos (-68.68) (rejected because cot x< 0).